How many particles are present in a closed container if the energy it contains is 95483.75J, and the diatomic oxygen gas is moving at a velocity of 48.11m/s? Use only the whole number for the value of atomic mass unit. Express your answer in proper scientific notation.
1
Expert's answer
2020-11-25T10:46:25-0500

Solution

Energy and number of particle relation is given by as

"E=\frac{nM_Av^2}{2N_A}"

"n=\frac{2E N_A}{M_Av^2}"


"n=\frac{2\times95483.75\times 6.023\times 10^{23}}{32\times(48.11)^2}"


"n=15.53\times10^{23}"